# Rapid Parameter Estimation for Merging Massive Black Hole Binaries Using ODE-Based Generative Models

### v1

@@ Abstract
Detecting the coalescences of massive black hole binaries (MBHBs) is one of the primary targets for space-based gravitational wave observatories such as LISA, Taiji, and Tianqin. The fast and accurate parameter estimation of merging MBHBs is of great significance for both astrophysics and the global fitting of all resolvable sources. However, such analyses entail significant computational costs. To address these challenges, inspired by the latest progress in generative models, we proposed a novel artificial intelligence (AI) based parameter estimation method called Variance Preserving Flow Matching Posterior Estimation (VPFMPE). Specifically, we utilize triangular interpolation to maintain variance over time, thereby constructing a transport path for training continuous normalization flows. Compared to the simple linear interpolation method used in flow matching to construct the optimal transport path, our approach better captures continuous temporal variations, making it more suitable for the parameter estimation of MBHBs. We also constructed a novel network architecture for gravitational wave feature extraction that outperforms existing methods in terms of training speed and accuracy. Additionally, we creatively introduce a parameter transformation method based on the symmetry in the detector’s response function. This transformation is integrated within VPFMPE, allowing us to train the model using a simplified dataset, and then perform parameter estimation on more general data, hence also acting as a crucial factor in improving the training speed. In conclusion, for the first time, within a comprehensive and reasonable parameter range, we have achieved a complete and unbiased 11-dimensional rapid inference for MBHBs in the presence of astrophysical confusion noise using ODE-based generative models. In the experiments based on simulated data, our model produces posterior distributions comparable to those obtained by nested sampling.
@@ 1 Introduction
When discussing the efficiency and speed of parameter estimation, the advancement in machine learning technology is revolutionizing our understanding of this area. Currently, most machine learning methods applied to the merging of MBHB use neural posterior estimation(NPE) [25, 26, 27] of discrete normalized flows. NPE has been widely used in fields such as atmospheric retrievals (AR) [28] and ground-based gravitational wave parameters estimation [29, 30]. NPE was also used by the winning entry to the 2023 edition of the ARIEL data challenge [31].
The mission lifetimes of LISA, Taiji, and Tianqin are planned to last for years, during which the payloads, orbits, and noise characteristics will vary over time. Accordingly, the models used for parameter estimation may also need to be regularly updated. As a result, the training speed of the model will be linked to the timeliness of scientific output. However, a significant challenge to the current NPE model in MBHB inference is the extensive training duration, which hampers its ability to conduct real-time data analysis. Additionally, previous research [32] in this area shows that NPE falls short in accurately reproducing the posterior distribution obtained by the benchmark Nested Sampling method. This prolonged training period not only delays the application of the model but also limits its scientific utility in rapidly evolving research contexts. To address these challenges, we applied the latest generative model to MBHBs and made further improvements and optimizations to better solve the problem. Compared to NPE, Flow Matching Posterior Estimation (FMPE) offers more flexibility in designing non-diffusion paths, such as optimal transport paths, and enables direct access to density functions [33]. Moreover, this method can achieve faster and more accurate results than NPE. Nevertheless, FMPE relies solely on simple linear interpolants when constructing optimal transport paths. This reliance impacts the extraction and modeling of MBHB waveform features, which in turn affects the efficiency and accuracy of parameter estimation.
Therefore, to capture the continuous temporal variations and enhance the parameter estimation of MBHBs, we aim to explore more effective methods for constructing transport paths and MBHB wave source feature extraction, which are particularly beneficial and targeted for extracting the continuous waveform characteristics of MBHBs. We consider utilizing trig interpolants to design transport paths [34, 35] that allow for the continuous evolution of MBHB waveform features along the time axis for modeling. This modeling accurately bridges any two probability density functions within a finite time, which belongs to a broad category of continuous-time stochastic processes. We call this method Variance Preserving Flow Matching Posterior Estimation (VPFMPE). Compared to standard Monte Carlo methods, the VPFMPE model achieves comparable performance.
@@
The challenges of parameter estimation for MBHBs not only involve the need for sophisticated algorithms capable of advanced feature extraction and inference, but also lie in the excessively wide prior range of parameters and the resulting complexity of data. Recent studies [36, 32] adopted priors that do not fully cover the complete physical ranges of parameters. This compromise was due to that setting broad prior ranges during model training can reduce the effectiveness of the model. This underscores the importance of finding a careful balance between the breadth of prior ranges and the accuracy of estimation. To date, no research has managed to achieve unbiased parameter estimation across the entire range of physical priors. Our method is proven to overcome this limitation. In this paper, we expanded the prior ranges of parameters to improve the applicability model.
@@
In order to further enhance the efficiency and effectiveness of model training, we tackled a common complexity faced by space-based gravitational wave detectors, namely the time-varying response function. Utilizing the coordinate transformation of parameters, we can essentially conduct both the model training and the parameter estimation in a coordinate frame fixed to the detector, in which the response function is time-invariant. This strategy significantly simplifies the training task. Consequently, the integration of coordinate transformation and advanced generative models enhances the speed and accuracy of parameter estimation for MBHBs across an extensive range of priors.
@@
In addition, we have developed a novel network architecture for gravitational wave feature extraction. This architecture illustrated in Figure 4, named TResMlp, introduces the innovation of encoding time $t$, resulting in superior training speed and accuracy compared to existing methods [33, 32, 36].
The training of the model can be finished in about two days, as shown in Figure 1, which is roughly three times faster than the training duration in our previous work [32]. Additionally, its performance on space-based gravitational wave data is on par with that of traditional Monte Carlo methods. This demonstrates the significant potential of integrating artificial intelligence with VPFMPE techniques to achieve results that are comparable to those obtained with standard Monte Carlo methods.
2 Data Generation and Preprocessing
In this section, we briefly describe the simulation of data used in model training and validation. The IMRPhenomD model is adopted to describe the source-frame waveforms of MBHBs. Given the widely recognized necessity of using Time Delay Interferometry (TDI) to suppress laser frequency noise, Taiji’s response to the MBHB signals is calculated in the form of two noise-orthogonal TDI channels $(A,E)$. Specifically, the fast Fourier domain transfer function derived by Ref. [37] is employed to project the IMRPhenomD waveform to the $(A,E)$ channels, which can be written in a concise form as
$\tilde{A}(f)=\sum_{\alpha}\mathcal{T}^{A}_{\alpha}(f,t_{f})\tilde{h}_{\alpha}(f),\quad\tilde{E}(f)=\sum_{\alpha}\mathcal{T}^{E}_{\alpha}(f,t_{f})\tilde{h}_{\alpha}(f),\quad\alpha\in\{+,\times\}.$ (1)
As can be seen, the transfer function has a close dependence on the orbits of the detector (through the $t_{f}$ term). For the sake of speed and simplicity, the motions of Taiji’s satellites are modeled based on an equal-arm analytic orbit with nominal arm-length $L=3\times 10^{9}{\rm m}$ and an orbital period of 1 year.
@@ 2 Data Generation and Preprocessing
The TDI response of each MBHB signal can be fully specified by 11 parameters: the chirp mass $\mathcal{M}_{c}$, mass ratio $q\equiv m_{2}/m_{1}$, the dimensionless spins along the $z-$axis $\chi_{z1}$ and $\chi_{z2}$, the time $t_{c}$ and phase $\varphi_{c}$ at coalescence, the luminosity distance $d_{L}$, the inclination angle $\iota$, polarization angle $\psi$, and the ecliptic coordinate $\{\lambda,\beta\}$, being the longitude and latitude respectively. The priors on these parameters are listed in Table 1. Similar to the LDC-sangria data, the prior range of $\mathcal{M}_{c}$ covers an order of magnitude. In our previous work [32], considering the complexity of data, we limited the prior ranges of spins to positive values. On this simplified parameter estimation task, NPE has been demonstrated to achieve unbiased estimations. However, its performance is subpar when applied to MBHB data with a complete spin prior. An advantage of the models proposed in this paper is that they are applicable to the complete physical prior range of spins (i.e. $[-1,1]$). Parameter $t_{c}$ is defined relative to a reference time $t_{\rm ref}$. Since the mergers of MBHBs manifest as prominent peaks in the time domain, we consider a prior range of $\sim 2000$ s ( $\sim 0.02$ day) around $t_{\rm ref}$. Notably, to simplify the dataset and reduce the difficulty of training, $t_{\rm ref}$ is fixed to the $1^{\rm th}$ day in the training set. The response function varies with the motion of detectors with a 1-year period, thus a training set with $t_{c}\in 1\pm 0.01$ day is far from being sufficient to cover all the observable MBHBs. While, thanks to the method described in Sec. 3.2, our model can be generalized to make inferences throughout the entire task cycle. Besides, the range of $d_{L}$ corresponds to the redshift range of $z\in[1,10]$, which is sufficient to encompass the majority of MBHBs observable by space-based GW detectors. The angles $\iota$ and $\beta$ are uniformly distributed in terms of $\cos\iota$ and $\sin\beta$ to represent that the sources are isotropically distributed on the sphere.
Table 1: The prior ranges of parameters. The last column is the alias of parameters used in the code as well as the corner plot.
@@
Parameter Description Prior Lower Bound Prior Upper Bound Alias in Code $\mathcal{M}_{c}$ Chirp mass of the binary system $1.0\times 10^{5}M_{\odot}$ $1.0\times 10^{6}M_{\odot}$ chirp_mass $q$ Ratio of masses of the two black holes 0.1 1.0 mass_ratio $\chi_{z1}$ Spin of the 1st black hole along the $z$-axis -1 1 spin_1 $\chi_{z2}$ Spin of the 2nd black hole along the $z$-axis -1 1 spin_2 $t_{c}$0.01 Time of coalescence relative to the reference time  [^1] The reference time $t_{\rm ref}$ is set to the $1^{\rm th}$ day., $t_{\rm ref}$ day 0.01 day coalescence_time $\varphi_{c}$ Phase at the moment of coalescence 0 rad $2\pi$ rad coalescence_phase $d_{L}$ Luminosity distance to the binary system 6000 Mpc 100000 Mpc luminosity_distance $\iota$ Angle of inclination of the binary orbit 0 rad $\pi$ rad inclination $\beta$ Ecliptic latitude of the binary system $-\frac{\pi}{2}$ rad $\frac{\pi}{2}$ rad latitude $\lambda$ Ecliptic longitude of the binary system 0 rad $2\pi$ rad longitude $\psi$ Polarization angle of the gravitational wave 0 rad $\pi$ rad psi
The majority of MBHB’s signal-to-noise ratio(SNR) is contributed by a short duration including the merger phase (typically several hours) [38]. To reasonably avoid the potential overlap of other signals, the time duration for each sample is set to one day, with the merger positioned towards the end of this window. Furthermore, the sampling frequency is set as $1/15$ Hz, indicating a Nyquist frequency of $0.033$ Hz. This frequency exceeds the highest instantaneous frequency of the waveforms, ensuring comprehensive capture of the signal.
@@
Figure 2: A representative data in the test set. The orange and green curves represent the Fourier transforms of the TDI-A, E data streams, which include GW signals, instrumental noises, and confusion noise. To clearly illustrate the relative strengths of the signal and noises, the blue curve depicts the Fourier transform of the confusion noise, while the gray dashed curve indicates the theoretical amplitude of the instrumental noises.
@@
Along with the detector responses, we also need the PSD of the noise in each channel. For instrumental noises, according to the current design of the Taiji mission, the one-sided ASDs of the two top-level instrumental noises, namely optical measurement system noise and test mass acceleration noise, are  [39]
@@
The total instrumental noise PSDs of the $A,E$ channels $S_{n}^{A,E}(f)$ are the combinations of these two components (see Ref. [40, 41, 42] for their specific forms). For simplicity, we treat $S_{n}(f)$ as constant over the whole mission duration. Our training data is a combination of “clean” TDI responses and stationary instrumental noises. On a more realistic ground, the long-term variations and short-term jitters of instrumental noise, along with the typical anomalies such as glitches and gaps should be considered in future work.
To test the robustness and generalizability of the model, during the inference stage, we consider a non-stationary noise caused by the unresolvable foreground originating from the overlapping GWs of GBs. This confusion noise is generated with the FastGB algorithm [43] from a catalog of $\sim 3\times 10^{7}$ GBs provided by the LDC dataset, from which $\sim 2\times 10^{4}$ bright sources with SNR greater than 7 were subtracted. This choice of SNR threshold follows the convention of [44, 45, 46, 47]. Shown in Figure 2 is representative data in the test set, including GW signal, instrumental noises, and confusion noise.
@@ 3 Methodological Framework
The application of machine learning techniques in the field of gravitational waves has seen widespread success and remarkable achievements [48, 49]. Among these techniques, deep learning has demonstrated exceptional prowess, especially in processing ground-based gravitational wave data [50, 51, 52, 53, 54, 55, 56, 57, 58].
NPE with normalizing flows. Neural Posterior Estimation (NPE) combined with Normalizing Flows (NF) is used to directly fit the posterior distribution $p(\theta|x)$. In this approach, an invertible mapping $\psi_{x}$ transforms the base distribution $p_{0}(\mathbf{z})$ into the target distribution $q(\theta|x)$:
@@
Continuous Normalizing Flows for parameter estimation. Continuous Normalizing Flows (CNFs) transform a simple base distribution into a more complex distribution by describing this transformation through continuous time $t\in[0,1]$. For each $t$, the flow is defined by a vector field $v_{t,x}$ on the sample space. This vector field represents the trajectory velocity of the samples at any given time $t$.
@@
The conversion between the base distribution (at $t=0$) and the target distribution (at $t=1$) is achieved by integration. A straightforward approach to training Continuous Normalizing Flows (CNFs) involves starting with a given base distribution and using an ODE solver to obtain the target distribution. This target distribution is then constrained to match the real data distribution by minimizing a divergence measure, such as the KL divergence.
Because many intermediate trajectories are unknown, inferring the distribution (either through sampling or likelihood computation) requires repeatedly simulating the ODE, leading to a significant computational burden.
However, Flow matching provides alternative training objectives for CNFs, Dax et al. [33] apply flow matching [59] to simulation-based inference. The loss function of FMPE is expressed as:
@@
Here, $v_{t,x}(\theta_{t})$ represents the vector field that generates the path to the target probability distribution, and $u_{t}(\theta_{t}|\theta_{1})$ is the sample-conditioned vector field. FMPE introduces a Gaussian path with time-dependent means $t\theta_{1}$ and standard deviations $1-(1-\sigma)t$.
@@
In our work, we reference the studies by Albergo et al. [34, 35] and utilize triangular interpolation to define the time-dependent mean and standard deviation, thus constructing the Gaussian path and sample-conditioned vector field as follows:
$p_{t}(\theta|\theta_{1})=N(\sin(\frac{\pi}{2t})\theta_{1}+\cos(\frac{\pi}{2t})\theta_{0},\sigma).$ (10)
$u_{t}(\theta|\theta_{1})=\frac{\pi}{2(\cos(\frac{\pi}{2t})\theta_{1}-\sin(\frac{\pi}{2t})\theta_{0})}.$ (11)
We call this method Variance Preserving Flow Matching Posterior Estimation and use it to train continuous normalization flows. We show the differences between the two methods in Table  2.
@@
VPFMPE trig $\sin(\frac{\pi}{2t})\theta_{1}+\cos(\frac{\pi}{2t})\theta_{0}$ $\sigma$
Figure 3: Comparison of the posterior distributions of the two methods. Blue represents the posterior distribution of the VPFMPE method on the injected MBHB signal. Orange represents the posterior distribution of the FMPE method on the injected MBHB signal. The red line represents the true value of the injected MBHB signal.
Comparison of methods. To demonstrate the performance of the transport path constructed by VPFMPE, we randomly injected an MBHB signal and compared the FMPE method with the VPFMPE method. Both models were trained using the same learning rate, optimizer, and gravitational wave feature extractor to ensure fairness. We show in Figure 3 a comparison of the posterior estimates of the four parameters of the injected MBHB signal by the two methods.
Discussion. The NPE network just only one-third allocated to the feature extractor network for GW data. However, the ODE-based generative model only needs to parameterize a simple vector field, so it can devote its full network capacity to interpreting the higher dimensions [33].

### v3

# Rapid Parameter Estimation for Merging Massive Black Hole Binaries Using Continuous Normalizing Flows

@@ Abstract
Detecting the coalescences of massive black hole binaries (MBHBs) is one of the primary targets for space-based gravitational wave observatories such as LISA, Taiji, and Tianqin. The fast and accurate parameter estimation of merging MBHBs is of great significance for the global fitting of all resolvable sources, as well as the astrophysical interpretation of gravitational wave signals. However, such analyses usually entail significant computational costs. To address these challenges, inspired by the latest progress in generative models, we explore the application of continuous normalizing flows (CNFs) on the parameter estimation of MBHBs. Specifically, we employ linear interpolation and trig interpolation methods to construct transport paths for training CNFs. Additionally, we creatively introduce a parameter transformation method based on the symmetry in the detector’s response function. This transformation is integrated within CNFs, allowing us to train the model using a simplified dataset, and then perform parameter estimation on more general data, hence also acting as a crucial factor in improving the training speed. In conclusion, for the first time, within a comprehensive and reasonable parameter range, we have achieved a complete and unbiased 11-dimensional rapid inference for MBHBs in the presence of astrophysical confusion noise using CNFs. In the experiments based on simulated data, our model produces posterior distributions comparable to those obtained by nested sampling.
@@ 1 Introduction
When discussing the efficiency and speed of parameter estimation, the advancement in machine learning technology is revolutionizing our understanding of this area. Currently, most machine learning methods applied to the merging of MBHB use neural posterior estimation (NPE) [25, 26, 27] of discrete normalized flows (DNFs). NPE has been widely used in fields such as atmospheric retrievals (AR) [28] and ground-based gravitational wave parameters estimation [29, 30]. NPE was also used by the winning entry to the 2023 edition of the ARIEL data challenge [31].
The mission lifetimes of LISA, Taiji, and Tianqin are planned to last for years, during which the payloads, orbits, and noise characteristics will vary over time. Accordingly, the models used for parameter estimation may also need to be regularly updated. As a result, the training speed of the model will be linked to the timeliness of scientific output. However, a significant challenge to the current NPE model in MBHB inference is the extensive training duration, which hampers its ability to conduct real-time data analysis. Additionally, previous research [32] in this area shows that NPE falls short in accurately reproducing the posterior distribution obtained by the benchmark Nested Sampling method. This prolonged training period not only delays the application of the model but also limits its scientific utility in rapidly evolving research contexts.
To address these challenges, we first apply Flow Matching (FM) [33] with Continuous Normalizing Flows (CNFs) to MBHBs. FM, as a training objective for CNFs, offers more flexibility in designing non-diffusion paths, such as optimal transport paths, and enables direct access to density functions [34]. This method is called Flow Matching Posterior Estimation (FMPE). The basic implementation of FM [33] uses simple linear interpolants to construct optimal transport paths. In our work, We also utilize trig interpolants [35, 36] to design transport paths that allow for the continuous evolution of MBHB waveform features along the time axis for modeling. We call this method Variance Preserving Flow Matching Posterior Estimation (VPFMPE).
@@
The challenges of parameter estimation for MBHBs not only involve the need for sophisticated algorithms capable of feature extraction and inference, but also lie in the excessively wide prior range of parameters and the resulting complexity of data. Recent studies [37, 32] adopted priors that do not fully cover the complete physical ranges of parameters. This compromise was due to that setting broad prior ranges during model training can reduce the effectiveness of the model. This underscores the importance of finding a careful balance between the breadth of prior ranges and the accuracy of estimation. To date, parameter estimation based on the full prior range of MBHB still remains to be explored. While, our method is proven to overcome this limitation and achieve unbiased parameter estimation on the full prior range, hence improving the applicability of AI models.
@@
In order to further enhance the efficiency and effectiveness of model training, we tackled a complexity commonly faced by space-based gravitational wave data analysis, namely the time-varying response function. Utilizing the coordinate transformation of parameters, we can essentially conduct both the model training and the parameter estimation in a coordinate frame fixed to the detector, in which the response function is time-invariant. This strategy significantly simplifies the training task. Consequently, the integration of coordinate transformation and advanced generative models enhances the speed and accuracy of parameter estimation for MBHBs.
@@
At last but not least, the network architecture developed in this work, named TResMlp, for GW feature extractions is illustrated in Fig 3, which introduces the innovation of encoding time $t$, resulting in superior training speed and accuracy compared to existing methods [34, 32, 37]. The training of the model can be finished in about two days, as shown in Figure 1, which is roughly three times faster than the training duration in previous work [32].
In conclusion, we performed a complete unbiased 11-dimensional fast inference of MBHB mergers using FM with CNFs. Compared to traditional Monte Carlo methods, we are at least several orders of magnitude faster, while still producing similar posterior distributions.
2 Data Generation
In this section, we introduce the simulation of datasets used for model training and testing. This part of work is a development based on our previous methods [32], with the most noteworthy improvement being the extension of prior ranges.
For the waveform template, we choose the IMRPhenomD template which describes the dominant $(\ell,|m|)=(2,2)$ harmonic of inspiral, merger and ringdown waveform for non-precessing binary black holes [38, 39]. To suppress laser frequency noise which dominates the raw interferometric data, a pre-processing technique known as Time Delay Interferometry (TDI) was proposed [40, 41], therefore we calculate Taiji’s response to MBHB signals in the form of two noise-orthogonal TDI channels $(A,E)$ [42]. The fast Fourier domain response function derived by Ref. [43] is employed to project the IMRPhenomD waveform to the $(A,E)$ TDI channels. The TDI response function can be written in a concise form as
$\tilde{A}(f)=\sum_{\alpha}\mathcal{T}^{A}_{\alpha}(f,t_{f})\tilde{h}_{\alpha}(f),\quad\tilde{E}(f)=\sum_{\alpha}\mathcal{T}^{E}_{\alpha}(f,t_{f})\tilde{h}_{\alpha}(f),$ (1)
where $\alpha$ denotes the polarization mode of waveform ($\alpha\in\{+,\times\}$). $\mathcal{T}_{\alpha}^{A,E}$ depend on the orbital motion of detector via the time-frequency relationship $t_{f}$
$t_{f}=-\frac{1}{2\pi}\frac{{\rm d}\Psi(f)}{{\rm d}f},$ (2)
with $\Psi(f)$ being the Fourier domain phase of the IMRPhenomD waveform. The motion of detector is modeled based on an equal-arm analytic orbit with nominal arm-length of $3\times 10^{9}\ {\rm m}$ and orbital period of 1 year. The Frequency domain TDI responses are calculated using a modified version of the open-source code BBHx [44, 45].
@@ 2 Data Generation
For the model considered in this paper, the parameter space of MBHB is 11-dimensional, whose symbols as well as descriptions can be found in Table 1. Also listed in Table 1 are the prior ranges of parameters. The explanations for the settings of priors are as follows. Similar to the LDC-sangria data, the prior range of $\mathcal{M}_{c}$ covers an order of magnitude. In our previous work [32], to reduce the complexity of data, we limited the prior ranges of spins to $\chi_{z1,2}\in[0,1]$, and NPE was demonstrated to achieve unbiased estimations on this simplified parameter estimation task. However, the performance of NPE became subpar when applied to the complete spin priors (i.e. $\chi_{z1,2}\in[-1,1]$). One of the key advantages of the models proposed in this paper is that they are applicable to the complete spin priors. Since the merger stage of each MBHB appears as a prominent peak in the time domain, we set the prior range of $t_{c}$ to be within $\sim 2000$ s ($\approx$ 0.02 day), with the central value denoted as the “reference time” $t_{\rm ref}$. Notably, to simplify the dataset and reduce the difficulty of training, in the training set, $t_{\rm ref}$ is fixed to the $1^{\rm th}$ day. The response function varies with the motion of detector with a 1-year period, thus a training set with $t_{c}\in[1-0.01,1+0.01]$ day is far from being sufficient to represent all the observable MBHBs. While, thanks to the method described in Sec. 3.2, our model can be generalized to make inferences on MBHB mergers throughout the entire mission lifetime. Consequently, when generating the test set, $t_{\rm ref}$ can take any value within the mission time. For other extrinsic parameters, the range of $d_{L}$ corresponds to the redshift range of $z\in[1,10]$, and the angles $\iota$ and $\beta$ are uniformly distributed in terms of $\cos\iota$ and $\sin\beta$, since the sources are isotropically oriented and located.
Table 1: The prior ranges of parameters. The table shows the range of prior distributions for each parameter.
@@
Parameter Description Prior Lower Bound Prior Upper Bound $\mathcal{M}_{c}$ Chirp mass of the binary system $1.0\times 10^{5}M_{\odot}$ $1.0\times 10^{6}M_{\odot}$ $q$ Ratio of masses of the two black holes 0.1 1.0 $\chi_{z1}$ Spin of the 1st black hole along the $z$-axis -1 1 $\chi_{z2}$ Spin of the 2nd black hole along the $z$-axis -1 1 $t_{c}$ Time of coalescence relative to the reference time -0.01 day 0.01 day $\varphi_{c}$ Phase at the moment of coalescence 0 rad $2\pi$ rad $d_{L}$ Luminosity distance to the binary system 6000 Mpc 100000 Mpc $\iota$ Angle of inclination of the binary orbit 0 rad $\pi$ rad $\beta$ Ecliptic latitude of the binary system $-\frac{\pi}{2}$ rad $\frac{\pi}{2}$ rad $\lambda$ Ecliptic longitude of the binary system 0 rad $2\pi$ rad $\psi$ Polarization angle of the gravitational wave 0 rad $\pi$ rad
The majority of a MBHB’s signal-to-noise ratio (SNR) is contributed by a short duration containing the merger stage (typically several hours) [46], therefore the time duration for each sample is set to one day. Besides, we set the sampling frequency of data to $1/15$ Hz, and we have checked that a Nyquist frequency of 0.033 Hz is beyond the maximum frequencies of all waveforms within the priors.
@@
Figure 2: A representative sample in the test set. The orange and green curves represent the Fourier transforms of the TDI-A, E data streams, which include GW signals, instrumental noises, and confusion noise. To clearly illustrate the relative strengths of the signal and noises, the blue curve depicts the Fourier transform of the confusion noise, while the gray dashed curve indicates the theoretical amplitude of the instrumental noises.
@@
Each sample in the training set is the combination of signal and instrumental noises. According to the current requirement of Taiji, the one-sided amplitude spectral densities (ASDs) of the two main instrumental noises, namely optical metrology system noise and test-mass acceleration noise, are  [47]
@@
and the ASDs of total instrumental noises of the $A,E$ channels (dubbed $\sqrt{S_{n}^{A,E}(f)}$) are the combinations of $S_{\mathrm{OMS}}(f)$ and $S_{\mathrm{ACC}}(f)$ (see Ref. [48, 49, 50] for their specific forms). We not only incorporated simulated noise into the MBHB signals during the training process but also utilized Taiji’s ASD function as additional contextual information fed to the model.
As is mentioned before, one of the advantages of our model is its robustness to the confusion noise. In the test set, we further include the foreground originating from the unresolvable GW signals of $\sim 3\times 10^{7}$ GBs. The simulation of confusion noise follows the methods of Refs. [51, 52, 53, 54, 55].
Shown in Figure 2 is a representative sample in the test set including GW signal, instrumental noises, and confusion noise.
@@ 3 Methodological Framework
The application of machine learning techniques in the field of gravitational wave data analysis has seen widespread success and remarkable achievements [56, 57]. Among these techniques, deep learning has demonstrated exceptional prowess, especially in processing the data of ground-based detectors [58, 59, 60, 61, 62, 63, 64, 65, 66]. Next, we will introduce the theoretical concepts of discrete and continuous flows, as well as the corresponding training objectives.
NPE with DNFs. NPE combined with DNFs is used to directly fit the posterior distribution $p(\theta|x)$. In this approach, an invertible mapping $\psi_{x}$ transforms the base distribution $p_{0}(\mathbf{z})$ into the target distribution $q(\theta|x)$:
@@
CNFs for parameter estimation. CNFs transform a simple base distribution into a more complex distribution by describing this transformation through continuous time $t\in[0,1]$. For each $t$, the flow is defined by a vector field $v_{t,x}$ on the sample space. This vector field represents the trajectory velocity of the samples at any given time $t$.
@@
The conversion between the base distribution (at $t=0$) and the target distribution (at $t=1$) is achieved by integration. A straightforward approach to training CNFs involves starting with a given base distribution and using an ODE solver to obtain the target distribution. This target distribution is then constrained to match the real data distribution by minimizing a divergence measure, such as the KL divergence. Because many intermediate trajectories are unknown, inferring the distribution (either through sampling or likelihood computation) requires repeatedly simulating the ODE, leading to a significant computational burden.
However, Flow matching provides a training objective for CNFs, Dax et al. [34] apply flow matching [33] to simulation-based inference. The loss function of FMPE is expressed as:
@@
Here, $v_{t,x}(\theta_{t})$ represents the vector field that generates the path to the target probability distribution, and $u_{t}(\theta_{t}|\theta_{1})$ is the sample-conditioned vector field. FMPE introduces a Gaussian path with time-dependent means $t\theta_{1}$ and standard deviations $1-(1-\sigma)t$. Overall, $t$ is sampled from $p(t)$ ($t\in[0,1]$), and the training data is generated by sampling $\theta_{1}$ from the prior and then simulating the corresponding data $x$ for $\theta_{1}$ [34]. The sample-conditional probability path must satisfy that at $t=0$, it is the base distribution (i.e., near $\theta_{0}$), and at $t=1$, it is the target distribution (i.e., near $\theta_{1}$).
@@
In our work, based on the studies by Albergo et al. [35, 36] and utilize triangular interpolation to define the time-dependent mean and standard deviation, thus constructing the Gaussian path and sample-conditioned vector field as follows:
$p_{t}(\theta|\theta_{1})=N(\sin(\frac{\pi t}{2})\theta_{1}+\cos(\frac{\pi t}{2})\theta_{0},\sigma).$ (11)
$u_{t}(\theta|\theta_{1})=\frac{\pi}{2(\cos(\frac{\pi t}{2})\theta_{1}-\sin(\frac{\pi t}{2})\theta_{0})}.$ (12)
We also use it to train continuous normalization flows. We show the differences between the two methods in Table  2.
@@
VPFMPE trig $\sin(\frac{\pi t}{2})\theta_{1}+\cos(\frac{\pi t}{2})\theta_{0}$ $\sigma$